OpenStudy (salemlover352):

Matt plans to put concrete on a rectangular portion of his driveway. The portion is 8 feet long and 4 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Matt needs is $58.07. Which of the following is closest to the width of the portion of the driveway on which Matt plans to put concrete?

OpenStudy (error1603):

Привет, 1 кубический ярд = 27 футов локоть, то мы должны найти No.of кубических ярдов .... 1 кубический ярд = $ 98 для "х" кубический ярд = $ 58.07 1 / х = 98 / 58.07 = 0,5925 х Теперь у нас есть преобразовать это и кубических футов ... 1 кубический ярд = 27 кубических футов 0,5925 кубических ярда = х кубических футов 1 / 0,5925 = 27 / хх = 15 .99 кубических футов в настоящее время ширина х высота х длина = 16 кубических футов WX 8 х (1/3) = 16 Вт = 48/8 = 6 футов да! там ширина двора примерно 6 футов .. надеюсь, что это будет я помочь!

OpenStudy (error1603):

Hi, 1 cubic yard = 27 cubit feet then we must find the no.of cubic yards .... for 1 cubic yard = $98 for "x" cubic yard = $58.07 1/x =98/58.07 x = 0.5925 Now we have to convert this in to cubic feet... 1 cubic yard = 27 cubic feet 0.5925 cubic yard = x cubic feet 1/0.5925= 27/x x = 15 .99 cubic feet now width x height x length = 16 cubic feet w x 8 x (1/3) = 16 w = 48/8 = 6 feet yep! there for the width of the yard is approximately 6 feet.. hope this will help ya !

OpenStudy (salemlover352):

Yes! Thank you!